# Likelihood-Scheduled Score-Based Generative Modeling for Fully 3D PET Image Reconstruction

### v1

@@ I Introduction
In this work, we propose a likelihood-scheduling mechanism for SGM-based reconstruction to address the aforementioned issues of burdensome hyperparameter tuning, slice inconsistency and slow reconstruction. Our method first runs the maximum likelihood expectation maximization algorithm (MLEM) to generate a “likelihood schedule” for a given set of measured sinogram data. The likelihood schedule is then integrated into the reverse diffusion process of an SGM-based reconstruction, enabling dynamic adjustment of the balance between the prior and the likelihood contributions. By integrating this development with SGMs trained on perpendicular slice orientations [16], our method eliminates the slice inconsistency issue while reducing the number of critical hyperparameters controlling the strength of regularization from 4 to 1.
Previous methods have implicitly altered the balance between likelihood and prior via regularization hyperparameters; our proposal is the first to investigate fixing the target likelihood upfront, providing samples from the posterior distribution of image reconstruction conditioned on both a likelihood value and noisy measured data.
@@
  • We compare methods for PET reconstruction with unsupervised SGMs against state-of-the-art conventional and supervised deep-learned reconstruction algorithms.
  • We resolve slice inconsistency issues in 3D by integrating our method with perpendicular pre-trained 2D SGMs.
  • We show and analyze the first-ever fully 3D PET reconstructions from real data with SGM-based reconstruction, specifically from low-count data acquired with radiotracer $[^{18}$]DPA714.
II Background
II-A PET reconstruction
@@
where $\mathbf{x}$ represents the true radiotracer distribution, $\mathbf{A}$ represents our system model and $\mathbf{b}$ models scatter and randoms components. The system model $\mathbf{A}$ accounts for positron range modeling, projection between image and sinogram space as well as attenuation and normalization modeling.
@@
However, with low-count data and a high-dimensional $\mathbf{x}$, the maximum likelihood estimate overfits to measurement noise. It is standard to compensate for this reduction in signal by conditioning on an image-based prior, thereby regularizing the reconstruction, via algorithms such as maximum a posteriori-expectation maximization (MAP-EM) [18]. For this purpose, let $p(\mathbf{x})$ be the prior probability density for an image $\mathbf{x}$.
@@
II-B Score-based generative models (SGMs)
@@
II-C PET reconstruction with SGMs
@@ III Related work
In addition to the above cited work on true reconstruction, some works have investigated the application of an SGM as a post-processing operator to PET image reconstruction, for the purpose of denoising [31, 32, 33, 34, 35].
@@ III-A Motivation
PET-DDS uses a modified DDIM sampling rule (see Section II-C), enforcing data consistency on Tweedie’s estimate via gradient ascent steps on a MAP proximal objective. This objective has three terms: the PLL $L_{j}$ for subset $j$, an anchor term to prevent straying too far from the diffusion output and an axial relative difference prior (RDP) for 3D reconstruction.
@@
IV Proposed approach
IV-A Likelihood-scheduling for SGM-based reconstruction
We propose to replace the fixed proximal data consistency update of existing methods with a dynamic data consistency update, that matches the likelihood of reconstruction iterates at each reverse diffusion step to a pre-computed ‘likelihood schedule’. For a visual explanation, see Fig. 1.
@@ IV-A Likelihood-scheduling for SGM-based reconstruction
The number of gradient ascent steps used in our algorithm is dynamic. The final reconstruction will have a PLL value within one gradient ascent step of a conventional MLEM reconstruction’s PLL value. This reconstruction approach may be viewed as sampling from the intersection of the prior distribution with a manifold of fixed likelihood images (see Fig. 2), as opposed to tuning hyperparameters that act as proxies for the likelihood-prior balance.
@@
Furthermore, it is easier to tune the critical hyperparameter $N_{\text{MLEM}}$, as it may be set to the number of EM updates for standard clinical reconstructions.
The proposed approach is more flexible to differing numbers of generative timesteps, as a likelihood schedule may be produced and adhered to for any number of generative timesteps. While hyperparameters such as the gradient ascent step size $\delta$ may still be specified, within reasonable bounds $\delta$ only controls the accuracy with which the likelihood schedule is conformed to, and not the balance of likelihood and prior.
IV-B Adaptations to 3D
@@
We instead take inspiration from the approach of Lee et al. [16], by leveraging pre-trained SGMs trained on slices from orthogonal orientations. Namely, we pre-train three SGMs $\mathbf{s}_{\mathbf{\theta}}^{\text{(sag)}},\mathbf{s}_{\mathbf{\theta}}^{\text{(cor)}},\mathbf{s}_{\mathbf{\theta}}^{\text{(tra)}}$ on diverse high-quality slices in the sagittal, coronal and transverse orientations respectively. During reconstruction, we apply each SGM to slices in its respective orientation and calculate the score $\mathbf{s}_{\mathbf{\theta}}(\mathbf{x}_{t},t)$ as the average of the score vectors output by the pre-trained SGMs. Note that this methodology differs from Lee et al. in that for each generative timestep, the sum of scores from 3 perpendicular SGMs is used rather than alternating the choice of score between 2 perpendicular SGMs.
@@
V Experimental setup
V-A Baseline methods
@@
V-A1 OSEM
As discussed in Section II-A, OSEM [19] is a model-based iterative method widely used on clinical scanners. In OSEM, expectation-maximization steps are taken with respect to subsets of the measured data, resulting in an accelerated version of MLEM. Regularization is implicitly achieved by early stopping before full convergence to the noisy maximum likelihood image estimate.
V-A2 MAP-EM
A more sophisticated conventional iterative method is MAP-EM [18], an iterative algorithm for maximizing a regularized Poisson likelihood function. For our implementation, we follow Wang & Qi’s formulation of patch-based edge-preserving regularization [36]. In this formulation, at each iteration an image estimate $\mathbf{x}^{OSEM}_{t+1}$ is computed via an OSEM update and a regularization image $\mathbf{x}^{reg}_{t+1}$ is calculated with Wang & Qi’s regularization function. The OSEM estimate and regularization image are then combined using the De Pierro update[37] (weighted by a scalar hyperparameter).
V-A3 FBSEM-net: unrolled iterative deep-learned method
@@
We compare to two implementations with different neural architectures for the regularizing neural network: FBSEM-net, with a convolutional neural network (CNN) comparable to Mehranian & Reader [5], and FBSEM-net-adv, with the same network architecture used for score-based learning (with a constant timestep input). Section V-E contains training details.
V-B PET forward operator
@@
The provided normalization was for data that has been axially compressed to span 11 (with 5 or 6 central lines of response summed along the axial direction). To cope with this constraint, axial compression was explicitly modeled in the forward operator (shown to have a negligible effect or reconstruction quality by Belzunce & Reader[39]). Positron range was approximated with a 4.5mm full-width half-maximum (FWHM) Gaussian point-spread function.
@@
where $\mathbf{x}$ is an image estimate, $N_{11}$ is a span 11 sinogram of normalization factors, $L_{11}$ is a span 11 sinogram of attenuation factors, $C_{11}$ is a compression operator that converts a span 1 sinogram into span 11, $X_{1}$ is the ParallelProj projector and $P$ is the Gaussian point-spread function.
V-C 3D real DPA-714 data
@@
The data had been previously acquired from 1-hour scans with the Siemens Biograph mMR, with approximately 200 MBq administered, with total counts in the range $2.9\times 10^{8}$ to $1.3\times 10^{9}$. At full-count, high-quality images (voxel size 2.08626 mm $\times$ 2.08626 mm $\times$ 2.03125 mm ; 3D image size $128\times 128\times 120$) were reconstructed with the scanner defaults (OSEM with 21 subsets, 2 iterations and no positron range modeling).
@@
V-D 2D simulated data
In order to have knowledge of the ground truth, simulated data were used. 2D transverse slices of high-quality DPA714 PET images were used as ground truths, with each image obtained via full-count reconstruction with Siemens’ implementation of the OSEM algorithm (with 21 subsets and 42 iterations). Forward projected data were obtained using a single axial ring, after rescaling each ground truth image slice such that the total count of simulated prompts matched the original estimate of true events. Corrective factors were modeled as purely attenuation and contamination was modeled as a constant background of 30% of simulated prompts.
@@
V-E Training and validation for SGMs and FBSEM-net
@@
V-F 2D reconstruction
Reconstructions in 2D were performed using simulated data (Section V-D). To calculate the 2D bias-variance assessment in Section VI-B and 2D likelihood-variance assessment in Section VI-F, for each of 10 random seeds, noisy sinogram data was generated according to the Poisson noise model. Then, reconstructions were performed for each independent noisy realization, with bias and standard deviation calculated according to Reader and Ellis[41].
@@
V-G 3D reconstruction
Reconstructions in 3D were performed from real clinical research data (see Section V-C). To match clinical software outputs, our ParallelProj forward operator was used without positron range modeling. To calculate the 3D likelihood-variance assessment in Section VI-G1, prompts and randoms were sampled at 10% of counts assuming independent Poisson statistics, with smoothed randoms and scatter sinograms re-estimated using Siemens’ scanner software. To accommodate the computational demands of 3D reconstruction, we perform only 25 diffusion steps per reconstruction, compute only a single reconstructed sample instead of a sample mean, and also use $\delta=1.0$.
@@
VI Results
VI-A 2D reconstruction performance
Table I shows the quantitative performance of each reconstruction method, assessed against 10 central 2D slices through each of 13 test datasets and averaged over 3 independent realizations of noisy sinogram data. Optimal hyperparameters for each method were established using a hyperparameter sweep to minimize NRMSE on the validation dataset.
The indicative reconstruction times listed in Table I allow us to conclude that the SGM-based methods are currently an order of magnitude slower than the conventional and supervised methods. For this set of optimal hyperparameters for PET-DDS, our proposed method was faster; this may not hold true in other settings.
TABLE I: Quantitative results for reconstruction methods applied to 2D DPA714 simulated data. 95% confidence intervals ($\pm$) calculated with respect to 3 different realizations of noisy measured data.
@@ VI Results
Figure 3: Reconstructions with each method from 2D simulated DPA714 data, with hyperparameters for each method chosen to minimize NRMSE on a validation dataset. Arrows point to key differences between reconstructions, including well-reconstructed structures (central arrow) and hallucinations due to the high noise level (right arrow). SGM-based reconstructions are the mean of 5 sampled reconstructions.
Figure 4: 2D bias-variance assessment for each method. For OSEM and MAP-EM, iteration number was varied. For PET-DDS, $\lambda$ was varied from 0 to 2, whereas for our method $N_{\text{MLEM}}$ was varied from 9 to 17.
VI-B 2D bias-variance assessment
Fig. 4 shows the results of a bias-variance assessment, performed on 10 central axial slices through the validation dataset. Where applicable, reconstruction hyperparameters were varied to show the effect of balancing the prior with the likelihood on the bias and variance properties of the reconstructions. This chart agrees with the previous quantitative results in Table I. In particular, our method achieves a similar or superior bias to PET-DDS with optimal hyperparameters for all variance levels.
VI-C 2D hyperparameter stability
To assess the stability of each SGM method, we considered varying the number of generative timesteps from 5 to 200 on both our method and PET-DDS, with the effect on reconstructions shown in Fig. 5. We also considered the effect of varying the step size $\delta$ of gradient ascent steps, shown in Fig. 6. Whereas PET-DDS fails to converge for $\delta\geq 0.8$, our method is robust to at least $\delta\leq 2.0$. Furthermore, our method with large $\delta$ uses fewer likelihood updates than generative steps, exhibiting remarkable efficiency.
@@
VI-D Sample path of likelihood-matched vs fixed updates

### v2

@@ 1 Introduction
In this work, we propose a likelihood-scheduling mechanism for SGM-based reconstruction to address the aforementioned issues of burdensome hyperparameter tuning, slice inconsistency and slow reconstruction. Our method first runs the maximum likelihood expectation maximization algorithm (MLEM) to generate a “likelihood schedule” for a given set of measured sinogram data (see Fig. 1). The likelihood schedule is then integrated into the reverse diffusion process of an SGM-based reconstruction, enabling dynamic adjustment of the balance between the prior and the likelihood contributions. By integrating this development with SGMs trained on perpendicular slice orientations [16], our method eliminates the slice inconsistency issue while reducing the number of critical regularization hyperparameters from 4 to 1.
Previous methods have implicitly altered the balance between likelihood and prior via regularization hyperparameters; our proposal is the first to investigate choosing the target likelihood upfront, providing samples from the posterior distribution of image reconstruction conditioned on both a likelihood value and noisy measured data.
@@
  • We resolve slice inconsistency issues in 3D by integrating our method with perpendicular pre-trained 2D SGMs and demonstrate the first-ever fully 3D PET reconstructions from real data using SGM-based reconstruction, specifically from low-count data acquired with radiotracer $[^{18}$]DPA714.
2 Background
2.1 PET reconstruction
@@
where $\mathbf{x}$ represents the true radiotracer distribution, $\mathbf{A}$ represents our system model and $\mathbf{b}$ models scatter and randoms components. The system model $\mathbf{A}$ includes the image space point spread function (PSF), projection between image and sinogram space as well as attenuation and normalization modeling.
@@
However, with low-count data and a high-dimensional $\mathbf{x}$, the maximum likelihood estimate overfits to noisy measurement data. It is standard to compensate for this reduction in signal by conditioning on an image-based prior, thereby regularizing the reconstruction, via algorithms such as maximum a posteriori-expectation maximization (MAP-EM) [18]. For this purpose, let $p(\mathbf{x})$ be the prior probability density for an image $\mathbf{x}$.
@@
2.2 Score-based generative models (SGMs)
@@
2.3 PET reconstruction with SGMs
@@ 3.1 Motivation
PET-DDS uses a modified DDIM sampling rule (see Section 2.3), enforcing data consistency on Tweedie’s estimate via gradient ascent steps on a MAP proximal objective. This objective has three terms: the PLL $L_{j}$ for subset $j$, an anchor term to prevent straying too far from the diffusion output and an axial relative difference prior (RDP) for 3D reconstruction.
@@
4 Proposed approach
4.1 Problem formulation
Suppose $D$ is the probability distribution over images learned by a pre-trained SGM. Let $c$ be a real scalar. Then, we seek to solve the following problem:
[…]
This problem formulation may be viewed as sampling from a manifold of fixed likelihood images (see Fig. 2), as weighted by their probability under the learned prior distribution.
For this problem to be meaningful, $c$ should be chosen such that clinically-relevant images exist with log-likelihood equal to $c$. In this work, we choose $c$ by computing a clinically-relevant image $\mathbf{x}_{\text{MLEM}}$ with the MLEM algorithm, and setting $c=L(\mathbf{x}_{\text{MLEM}}|\mathbf{m})$. This is a desirable selection, as solving the above problem would lead to sampled images that are equally consistent with measured sinograms as MLEM images, but without the issues of early-terminated MLEM (chiefly a lack of resolved detail).
4.2 Likelihood-scheduling for SGM-based reconstruction
We propose to solve the problem defined in Section 4.1 with a dynamic data consistency update that matches the likelihood of reconstruction iterates at each reverse diffusion step to a pre-computed ‘likelihood schedule’. For a visual explanation, see Fig. 1.
@@ 4.2 Likelihood-scheduling for SGM-based reconstruction
The number of gradient ascent steps used in our algorithm is dynamic. The final reconstruction will have a PLL value within one gradient ascent step of a conventional MLEM reconstruction’s PLL value.
@@
Furthermore, it is easier to tune the critical hyperparameter $N_{\text{MLEM}}$, as it may be set to the number of EM updates for standard clinical reconstructions. All unsupervised PET reconstruction approaches involve heuristic regularization hyperparameters; a particular strength of our approach is to align this heuristic with an existing clinically-accepted and vendor-recommended heuristic, i.e. $N_{\text{MLEM}}$.
The proposed approach is also more flexible to differing numbers of generative timesteps, as a likelihood schedule may be produced and adhered to for any number of generative timesteps. While hyperparameters such as the gradient ascent step size $\delta$ may still be specified, within reasonable bounds $\delta$ only controls the accuracy with which the likelihood schedule is conformed to, and not the balance of likelihood and prior. This is fundamentally different from gradient-ascent scheduling approaches such as linear annealing, which do not reduce the number of hyperparameters required.
4.3 Adaptations to 3D
@@
We instead take inspiration from the approach of Lee et al. [16], by leveraging pre-trained SGMs trained on slices from orthogonal orientations. Namely, we pre-train three SGMs $\mathbf{s}_{\mathbf{\theta}}^{\text{(sag)}},\mathbf{s}_{\mathbf{\theta}}^{\text{(cor)}},\mathbf{s}_{\mathbf{\theta}}^{\text{(tra)}}$ on diverse high-quality slices in the sagittal, coronal and transverse orientations respectively. During reconstruction, we apply each SGM to slices in its respective orientation and calculate the score $\mathbf{s}_{\mathbf{\theta}}(\mathbf{x}_{t},t)$ as the average of the score vectors output by the pre-trained SGMs. Note that this methodology differs from Lee et al. in that for each generative timestep, the sum of scores from 3 perpendicular SGMs is used rather than alternating the choice of score between 2 perpendicular SGMs. Alternating between directions leads to Tweedie’s estimate iterates with slice inconsistency effects, which are eliminates gradually as many diffusion steps are taken. Our different approach was hence motivated by the desire to reduce the number of diffusion timesteps (due to the expense of the likelihood updates) while eliminating slice inconsistency effects throughout the diffusion process.
@@
5 Experimental setup
5.1 Baseline methods
@@
5.1.1 OSEM
As discussed in Section 2.1, OSEM [19] is a model-based iterative method widely used on clinical scanners. In OSEM, expectation-maximization steps are taken with respect to subsets of the measured data, resulting in an accelerated version of MLEM. Regularization is implicitly achieved by early stopping before full convergence to the noisy maximum likelihood image estimate.
5.1.2 MAP-EM
A more sophisticated conventional iterative method is MAP-EM [18], an iterative algorithm for maximizing a regularized Poisson likelihood function. For our implementation, we follow Wang & Qi’s formulation of patch-based edge-preserving regularization [31]. In this formulation, at each iteration an image estimate $\mathbf{x}^{OSEM}_{t+1}$ is computed via an OSEM update and a regularization image $\mathbf{x}^{reg}_{t+1}$ is calculated with Wang & Qi’s regularization function. The OSEM estimate and regularization image are then combined using the De Pierro update[32] weighted by a scalar hyperparameter $\beta$.
5.1.3 FBSEM-net: unrolled iterative deep-learned method
@@
We compare to two implementations with different neural architectures for the regularizing neural network: FBSEM-net, with a convolutional neural network (CNN) comparable to Mehranian & Reader [5], and FBSEM-net-adv, with the same network architecture used for score-based learning (with a constant timestep input). Section 5.5 contains training details.
FBSEM-net-adv is included in this study as a strong baseline representing the performance of deep learning approaches on within-distribution datasets when sinogram training data is available. Despite its strong performance in 2D, FBSEM-net-adv is omitted in 3D, due to computational infeasibility to train on available hardware.
5.2 PET forward operator
@@
The provided normalization was for data that has been axially compressed to span 11 (with 5 or 6 central lines of response summed along the axial direction). To cope with this constraint, axial compression was explicitly modeled in the forward operator (shown to have a negligible effect on reconstruction quality by Belzunce & Reader[34]). In 2D, a 4.5mm full-width half-maximum (FWHM) Gaussian PSF was also used.
@@
where $\mathbf{x}$ is an image estimate, $N_{11}$ is a span 11 sinogram of normalization factors, $L_{11}$ is a span 11 sinogram of attenuation factors, $C_{11}$ is a compression operator that converts a span 1 sinogram into span 11, $X_{1}$ is the ParallelProj projector and $P$ is the Gaussian PSF.
5.3 3D real $[^{18}$]DPA-714 data
@@
The data had been previously acquired from 1-hour scans with the Siemens Biograph mMR, with approximately 200 MBq administered, with total counts in the range $2.9\times 10^{8}$ to $1.3\times 10^{9}$. At full-count, high-quality images (voxel size 2.08626 mm $\times$ 2.08626 mm $\times$ 2.03125 mm ; 3D image size $128\times 128\times 120$) were reconstructed with the scanner defaults (OSEM with 21 subsets, 2 iterations and no PSF).
@@
5.4 2D simulated data
In order to have knowledge of the ground truth, simulated data were used. 2D transverse slices of high-quality $[^{18}$]DPA-714 PET images were used as ground truths, with each image obtained via full-count reconstruction with Siemens’ implementation of the OSEM algorithm (with 21 subsets and 42 iterations, i.e. 2 full passes through the data). Forward projected data were obtained using a single axial ring, after rescaling each ground truth image slice such that the total count of simulated prompts matched the original estimate of true events. Corrective factors were modeled as purely attenuation and contamination was modeled as a constant background of 30% of simulated prompts.
@@
5.5 Training and validation for SGMs and FBSEM-net
@@
Figure 3: Reconstructions with each method from 2D simulated $[^{18}$]DPA-714 data, with hyperparameters for each method chosen to minimize NRMSE on a validation dataset. Arrows point to key differences between reconstructions, including well-reconstructed structures (central arrow) and hallucinations due to the high noise level (right arrow). SGM-based reconstructions are the mean of 5 sampled reconstructions.
5.6 2D reconstruction
Reconstructions in 2D were performed using simulated data (Section 5.4). To calculate the 2D bias-variance assessment in Section 6.2 and 2D likelihood-variance assessment in Section 6.7, for each of 10 random seeds, noisy sinogram data was generated according to the Poisson noise model. Then, reconstructions were performed for each independent noisy realization, with bias and standard deviation calculated according to Reader and Ellis[36].
@@
5.7 3D reconstruction
Reconstructions in 3D were performed from real clinical research data (see Section 5.3). To match clinical software outputs, our ParallelProj forward operator was used without a PSF. To calculate the 3D likelihood-variance assessment in Section 6.8.2, prompts and randoms were sampled at 10% of counts assuming independent Poisson statistics, with smoothed randoms and scatter sinograms re-estimated using Siemens’ scanner software. To accommodate the computational demands of 3D reconstruction, we perform only 25 diffusion steps per reconstruction, compute only a single reconstructed sample instead of a sample mean, and also use $\delta=1.0$.
@@
6 Results
6.1 2D reconstruction performance
Table 1 shows the quantitative performance of each reconstruction method, assessed against 10 central 2D slices through each of 13 test datasets and averaged over 3 independent realizations of noisy sinogram data. Optimal hyperparameters for each method were established using a hyperparameter sweep to minimize NRMSE on the validation dataset.
Table 1: Quantitative results for reconstruction methods applied to 2D $[^{18}$]DPA-714 simulated data, with hyperparameters for each method chosen to minimize NRMSE on a validation dataset. 95% confidence intervals ($\pm$) calculated with respect to 3 different realizations of noisy measured data.
@@ 6 Results
The indicative reconstruction times listed in Table 1 allow us to conclude that the SGM-based methods are currently an order of magnitude slower than the conventional and supervised methods. For this set of optimal hyperparameters for PET-DDS, our proposed method was faster; this may not hold true in other settings.
Figure 4: 2D bias-variance assessment for each method [36]. For OSEM and MAP-EM ($\beta=3$), one subset was used and iteration number was varied from 5 to 100. For PET-DDS, $\lambda$ was varied from 0 to 2 (with $\lambda=0$ the rightmost point and $\lambda$ values increasing in increments of 0.5 for each of the data points displayed), whereas for our method $N_{\text{MLEM}}$ was varied from 9 to 17 (increments of 2 shown, rightmost point corresponds to $N_{\text{MLEM}}=9$.
6.2 2D bias-variance assessment
Fig. 4 shows the results of a bias-variance assessment, performed on 10 central axial slices through the validation dataset. Where applicable, reconstruction hyperparameters were varied to show the effect of balancing the prior with the likelihood on the bias and variance properties of the reconstructions. This chart agrees with the previous quantitative results in Table 1. In particular, our method achieves a similar or superior bias to PET-DDS with optimal hyperparameters for all variance levels.
6.3 2D hyperparameter stability
To assess the stability of each SGM method, we considered varying the number of generative timesteps from 5 to 200 on both our method and PET-DDS, with the effect on reconstructions shown in Fig. 5. We also considered the effect of varying the step size $\delta$ of gradient ascent steps, shown in Fig. 6. Whereas PET-DDS fails to converge for $\delta\geq 0.8$, our method is robust to at least $\delta\leq 2.0$. Furthermore, our method with large $\delta$ uses fewer likelihood updates than generative steps, exhibiting remarkable efficiency, and demonstrating that our method is more efficiently exploiting the reduced need for gradient ascent steps with a strong diffusion prior.
@@
6.4 Sample path of likelihood-matched vs fixed updates